# Intrinsic Mixed-state Quantum Topological Order

### v2

@@ III Generalities: anomalous 1-form symmetry, nontrivial statistics, and long-range entanglement
The long-range entanglement of the gapless spin liquid and the intrinsic mixed-state quantum TO are both related to an anomalous 1-form symmetry, which is generated by the following loop operators
@@
Now we address another sublte question: What does “deconfinement” really mean in mixed states under noisy channels? Conventionally, it means that the energy cost does not grow indefinitely by separating individual topological excitations far apart. Here this definition does not make sense because the system is no longer governed by a Hamiltonian. Therefore, we propose the following definition of deconfined excitations for generic mixed states.
@@
Clearly, the above definition is consistent with the conventional notion of deconfined excitations, and serves as a faithful and natural generalization to open quantum systems. Next, we illustrate that $f$ anyons (we omit the “+” hereafter) are indeed deconfined excitations according to this definition, created by the unitary operator $U_{\tilde{C}}$. The second condition can be easily verified using the strong $1$-form symmetry. We denote the support of $O$ as $\Omega(O)$ for convenience, which we assume to be away from $i,j$. If $\Omega(O)\bigcap\tilde{C}=\emptyset$, then $\text{tr}(U_{\tilde{C}}\rho U_{\tilde{C}}^{\dagger}O)=\text{tr}(\rho U_{\tilde{C}}^{\dagger}OU_{\tilde{C}})=\text{tr}(\rho O)$; if $\Omega(O)\bigcap\tilde{C}\neq\emptyset$, we can always find another open string $\tilde{C}^{\prime}$, such that $\tilde{C}\bigcup\tilde{C}^{\prime}$ is a contractible loop and $\Omega(O)\bigcap\tilde{C}^{\prime}=\emptyset$. Using the strong $1$-form symmetry $U_{\tilde{C}^{\prime}}^{\dagger}U_{\tilde{C}}\rho_{f}=\rho_{f}$, it is straightforward to get $\text{tr}(U_{\tilde{C}}\rho U_{\tilde{C}}^{\dagger}O)=\text{tr}(U_{\tilde{C}^{\prime}}\rho U_{\tilde{C}^{\prime}}^{\dagger}O)=\text{tr}(\rho O)$. The first condition follows from the fermionic statistics of the $f$ anyons Levin and Wen (2003). See also the discussion below and in Section III.
@@
Is the fermionic statistics well-defined for mixed states? We give an affirmative answer with the following microscopic method of detection (see Fig. 3 for an illustration): $a.$ Create two $f$ anyons at locations $i$, $j$ using $W^{f}_{ji}=W^{f}_{jk}W^{f}_{ki}$; $b.$ move the $f$ anyon at $i$ to location $l$ using $W^{f}_{li}=W^{f}_{lk}W^{f}_{ki}$. $c.$ move the other $f$ anyon from $j$ to $i$ using $W^{f}_{ij}=W^{f\dagger}_{ji}$; $d.$ move the $f$ anyon at $l$ to location $j$ using $W^{f}_{jl}=W^{f}_{jk}W^{f}_{kl}$; $e.$ annihilate the two $f$ anyons using $W^{f}_{ij}$. In step $b$-$d$ the locations of the two $f$ anyons are exchanged, which result in a phase $W^{f}_{jl}W^{f}_{ij}W^{f}_{li}=\theta(f)=-1$. To turn this statistical phase into an observable effect, we can use the same protocol as in Satzinger et al. (2021). That is, we introduce an ancilla qubit and prepare the initial state $|+\rangle\langle+|\otimes\rho_{f}$, where $|+\rangle\equiv\frac{|1\rangle+|0\rangle}{\sqrt{2}}$. Then we can use the ancilla qubit to control the exchange process. Namely, we perform step $a$-$e$ when the ancilla is in state $|1\rangle$, and do nothing otherwise. This controlled process can be performed using the unitary operator $V=\frac{1}{\sqrt{2}}(|0\rangle\langle 0|\otimes I+|1\rangle\langle 1|\otimes W^{f}_{ij}W^{f}_{jl}W^{f}_{ij}W^{f}_{li}W^{f}_{ji})$. Then the statistical phase will manifest as the rotation of the ancilla:
@@
where $|-\rangle=\frac{|0\rangle-|1\rangle}{\sqrt{2}}$. The fermionic statistics of emergent deconfined excitations is the most striking observable effect of the intrinsic mixed-state quantum TO. Here, it is guaranteed by the anomalous strong $1$-form symmetry.
In the double space, $e_{+}e_{-}$ also remains a deconfined excitation, which has mutual semion statistics with $f_{+}$. Surprisingly, this braiding statistics can also be (partially) detected in the physical Hilbert space, as we demonstrate below. First, note that $\rho_{f}$ preserves a weak $1$-form symmetry generated by $e$ anyons, $W^{e}_{\gamma}\rho_{f}W^{e}_{\gamma}=\rho$ for contractible closed $e$ string $W^{e}_{\gamma}$ de Groot et al. (2022). Then based on Definition 1, deconfined excitations can be created using an open $e$ string, $\rho_{f}\rightarrow\rho^{e}_{f}=W^{e}_{C}\rho_{f}W^{e}_{C}$. Secondly, introduce an ancilla qubit and prepare the initial state $|+\rangle\langle+|\otimes\rho_{f}^{e}$. Then use the ancilla qubit to control the braiding process: if the ancilla is in $|1\rangle$, we create two $f$ anyons and drag one of them along a loop $l$ enclosing the $e$ anyon, and finally annihilate with the other $f$ anyon; if the ancilla is in $|0\rangle$ we do nothing. This controlled process can be performed using the unitary gate $V=\frac{1}{\sqrt{2}}(|0\rangle\langle 0|\otimes I+|1\rangle\langle 1|\otimes W^{f}_{\tilde{\gamma}}$, where $W^{f}_{\tilde{\gamma}}=\prod_{i\in\text{loop }\tilde{\gamma}}X_{i}Z_{i+\bm{\delta}}$. Then the braiding statistics can be detected by rotation of the ancilla qubit:
@@
One key distinction to the usual anyon braiding is that the $f-e$ braiding here is only one-way defined. Namely, the statistical phase can only be detected by moving $f$ around $e$ (in a way that is controlled by the ancilla, as demonstrated above), but not the other way around. This directly follows from the fact that while $f$ anyons generate a strong $1$-form symmetry, $e$ anyons only generate a weak $1$-form symmetry. Therefore, a finer characterization of deconfined excitations than Definition 1 is needed, and we call the former type as strongly deconfined and the latter type as weakly deconfined. As demonstrated above, strongly deconfined anyons are allowed to form coherent superposition states, and their statistics can be detected from the interference effects. Weakly deconfined anyons, on the other hand, always have trivial exchange and braiding statistics among themselves and can have one-way braiding statistics with strongly deconfined anyons. In the rest of the paper we mainly focus on strongly deconfined anyons.
@@
Using the perspective of anomalies, we can generalize our construction to obtain other intrinsic mixed-state quantum TO  [^2] This manuscript has been substantially revised since our first submission in July 2023. In the meantime, we became aware of two recent works Sohal and Prem (2024); Ellison and Cheng (2024) on the same topic. The discussion regarding the non-modular anyon theory in Section IV.2 is inspired by these two works, where a connection between the intrinsic mixed-state TO and the topological subsystem code is established Ellison et al. (2023). The perspective of $1$-form symmetry anomaly is also adopted in another recent paper written by one of us Wang and Li (2024).. Given a TO with anyon content $\mathcal{A}$, one can incoherently proliferate a subset $\mathcal{B}\subset\mathcal{A}$ of anyons using a noisy channel, taking the Kraus operator to be the shortest string of anyons in $\mathcal{B}$. If there exists some Abelian anyon $a\in\mathcal{A}$ that have trivial braiding with all anyons in $\mathcal{B}$, then closed $a$ strings commute with the Kraus operators. Therefore, the $a$ anyons generate a strong $1$-form symmetry. Based on the previous discussion, it means they remain deconfined excitations with well-defined statistics. When they have nontrivial statistics, $\theta(a)\neq 1$, then the 1-form symmetry is anomalous. For example, in the model discussed above, $\mathcal{A}=\{1,e,m,f\},\mathcal{B}=\{f\}$, and $a=f$ because fermions have trivial full braiding with themselves.
The existence of non-bosonic strongly deconfined excitation can be viewed as another diagnosis of nontrivial mixed-state quantum TO, in complementary to TEN. Indeed, we show below that this property generically implies the mixed state to be LRE, without explicit calculation of negativity.
@@
    Typically, the 1-form symmetry includes generators supported on non-contractible loops. However, here we only require the symmetry condition for contractible ones, enabling a much wider range of applications. For example, for the construction of intrinsic mixed-state quantum TO discussed above, the pre-decoherence state can be taken to be any state in the ground state subspace. The above theorem can also be applied to topologically trivial spatial manifold like a $2$-sphere.
@@
    To prove the theorem, we make an additional assumption that $W^{a}_{\gamma}$ can be expressed using a FDLUC, which is usually the case for Abelian anyon $a$.
Proof. Suppose by contradiction that $\rho=\mathcal{N}_{\text{FDLC}}[\sum_{\lambda}p_{\lambda}\bigotimes_{\text{site}i}|\psi^{\lambda}_{i}\rangle\langle\psi^{\lambda}_{i}|]$. Recall that a generic FDLC $\mathcal{N}$ can be constructed in three steps: First, introduce auxiliary degrees of freedom on each site. Secondly, apply a finite-depth local unitary circuit (FDLUC) on the system and the auxiliary degrees of freedom. Finally, trace out the added degrees of freedom. It can be written as $\mathcal{N}_{\text{FDLC}}[\rho_{0}]=\text{tr}_{E}[U_{SE}\rho_{0}\bigotimes_{\text{site i}}|e_{i}\rangle\langle e_{i}|U^{\dagger}_{SE}]$, where $U_{SE}$ is a FDLUC. If $\rho_{0}$ is a fully separable state, then $\rho=\mathcal{N}_{\text{FDLC}}[\rho_{0}]$ is of the form $\rho=\sum_{\lambda}p_{\lambda}\text{tr}_{E}[U_{SE}|\phi_{i}^{\lambda}\rangle\langle\phi_{i}^{\lambda}|U^{\dagger}_{SE}]$, where $|\phi_{i}\rangle\equiv|\psi_{i}^{\lambda}\rangle\otimes|e_{i}^{\lambda}\rangle$ Then each pure state $U_{SE}\bigotimes_{i}|{\phi}_{i}\rangle$ must be symmetric: $W^{a}_{\gamma}U_{SE}\bigotimes_{i}|\phi_{i}\rangle=U_{SE}\bigotimes_{i}|\phi_{i}\rangle$. Equivalently, $\tilde{W}^{a}_{\gamma}=U^{\dagger}_{SE}{W}_{\gamma}^{a}U_{SE}$ is a symmetry of the product state $\bigotimes_{i}|\phi_{i}\rangle$. Crucially, the restriction of $\tilde{W}^{a}$ to open strings $\tilde{W}^{a}_{i_{1}i_{2}}\equiv U^{\dagger}_{SE}{W}_{i_{1}i_{2}}^{a}U_{SE}$ satisfies the same algebra (18) as $W^{a}$. As we show below, this will lead to contradiction.
@@
By applying both sides of the equation on $\bigotimes_{i}|\tilde{\psi_{i}}\rangle$, we get $A_{k}\otimes B_{j}\otimes A_{l}\otimes B_{m}\bigotimes_{i}|\phi_{i}\rangle=\theta(a)A_{k}\otimes B_{m}\otimes A_{l}\otimes B_{j}\bigotimes_{i}|\phi_{j}\rangle$, which leads to contradiction when $\theta(a)\neq 1$.
The above proof shows that nontrivial quantum TO is guaranteed by the existence of non-bosonic deconfined excitation $a$, even for mixed states. Furthermore, when $a$ is neither bosonic or fermionic, $\theta(a)\neq\pm 1,$ or when $a$ has nontrivial mutual statistics with another strongly deconfined anyon $a^{\prime}$, then an even stronger conclusion can be proved.
Theorem 2: Consider a $2$-dimensional state $\rho$ with an anomalous strong 1-form symmetry generated by Abelian anyons $a,a^{\prime}$ with nontrivial braiding statistics, $B_{\theta}(a,a^{\prime})\neq 1$. Then for any bipartition $A\bigcup B$ with linear size $L_{A},L_{B}\rightarrow\infty$ in the thermodynamic limit, $\rho$ cannot be prepared using a FDLC from any bipartite separable state:
@@
$\tilde{W}^{a}_{j^{\prime}_{A}i^{\prime}_{A}}\tilde{W}^{a}_{j_{A}i_{A}}=B_{\theta}(a,a^{\prime})\tilde{W}^{a}_{j_{A}i_{A}}\tilde{W}^{a^{\prime}}_{j^{\prime}_{A}i^{\prime}_{A}}$ (25)
$\Rightarrow$ $u^{\dagger}_{\sigma_{i^{\prime}j^{\prime}}}\tilde{W}^{a}_{j^{\prime}_{A}i^{\prime}_{A}}u^{\dagger}_{\sigma_{ij}}\tilde{W}^{a}_{j_{A}i_{A}}=B_{\theta}(a,a^{\prime})u^{\dagger}_{\sigma_{ij}}\tilde{W}^{a}_{j_{A}i_{A}}u^{\dagger}_{\sigma_{i^{\prime}j^{\prime}}}\tilde{W}^{a}_{j^{\prime}_{A}i^{\prime}_{A}},$
@@
IV Generalizations to other intrinsic mixed-state quantum TO

### v3

@@ III Generalities: anomalous 1-form symmetry, nontrivial statistics, and long-range entanglement
The long-range entanglement of the gapless spin liquid and the intrinsic mixed-state TO are both related to an anomalous 1-form symmetry, which is generated by the following loop operators
@@
Now we address another subtle question: What does “deconfinement” really mean in mixed states under noisy channels? Conventionally, it means that the energy cost does not grow indefinitely by separating individual topological excitations far apart. Here this definition does not make sense because the system is no longer governed by a Hamiltonian. Therefore, we propose the following definition of deconfined excitations for generic mixed states.
@@
Clearly, the above definition is consistent with the conventional notion of deconfined excitations, and serves as a faithful and natural generalization to open quantum systems. Next, we illustrate that $f$ anyons (we omit the “+” hereafter) are indeed deconfined excitations according to this definition, created by the unitary operator $U_{\tilde{C}}$. The second condition can be easily verified using the strong $1$-form symmetry. We denote the support of $O$ as $\Omega(O)$ for convenience, which we assume to be away from $i,j$. If $\Omega(O)\bigcap\tilde{C}=\emptyset$, then $\text{tr}(U_{\tilde{C}}\rho U_{\tilde{C}}^{\dagger}O)=\text{tr}(\rho U_{\tilde{C}}^{\dagger}OU_{\tilde{C}})=\text{tr}(\rho O)$; if $\Omega(O)\bigcap\tilde{C}\neq\emptyset$, we can always find another open string $\tilde{C}^{\prime}$, such that $\tilde{C}\bigcup\tilde{C}^{\prime}$ is a contractible loop and $\Omega(O)\bigcap\tilde{C}^{\prime}=\emptyset$. Using the strong $1$-form symmetry $U_{\tilde{C}^{\prime}}^{\dagger}U_{\tilde{C}}\rho_{f}=\rho_{f}$, it is straightforward to get $\text{tr}(U_{\tilde{C}}\rho U_{\tilde{C}}^{\dagger}O)=\text{tr}(U_{\tilde{C}^{\prime}}\rho U_{\tilde{C}^{\prime}}^{\dagger}O)=\text{tr}(\rho O)$. The first condition follows from the fermionic statistics of the $f$ anyons Levin and Wen (2003). See also the proof of Theorem 1 for more details.
@@
Is the fermionic statistics well-defined for mixed states? We give an affirmative answer with the following microscopic method of detection (see Fig. 3 for an illustration): $a.$ Create two $f$ anyons at locations $i$, $j$ using $W^{f}_{ji}=W^{f}_{jk}W^{f}_{ki}$; $b.$ move the $f$ anyon at $i$ to location $l$ using $W^{f}_{li}=W^{f}_{lk}W^{f}_{ki}$. $c.$ move the other $f$ anyon from $j$ to $i$ using $W^{f}_{ij}=W^{f\dagger}_{ji}$; $d.$ move the $f$ anyon at $l$ to location $j$ using $W^{f}_{jl}=W^{f}_{jk}W^{f}_{kl}$; $e.$ annihilate the two $f$ anyons using $W^{f}_{ij}$. In step $b$-$d$ the locations of the two $f$ anyons are exchanged, which result in a phase $W^{f}_{jl}W^{f}_{ij}W^{f}_{li}=\theta(f)=-1$. To turn this statistical phase into an observable effect, we can use the same protocol as in Satzinger et al. (2021). That is, we introduce an ancilla qubit and prepare the initial state $|+\rangle\langle+|\otimes\rho_{f}$, where $|+\rangle\equiv\frac{|1\rangle+|0\rangle}{\sqrt{2}}$. Then we can use the ancilla qubit to control the exchange process. Namely, we perform step $a$-$e$ when the ancilla is in state $|1\rangle$, and do nothing otherwise. This controlled process can be performed using the unitary operator $V=|0\rangle\langle 0|\otimes I+|1\rangle\langle 1|\otimes W^{f}_{ij}W^{f}_{jl}W^{f}_{ij}W^{f}_{li}W^{f}_{ji}$. Then the statistical phase will manifest as the rotation of the ancilla:
@@
where $|-\rangle=\frac{|0\rangle-|1\rangle}{\sqrt{2}}$. The fermionic statistics of emergent deconfined excitations is the most striking observable effect of the intrinsic mixed-state TO. Here, it is guaranteed by the anomalous strong $1$-form symmetry.
In the double space, $e_{+}e_{-}$ also remains a deconfined excitation, which has mutual semion statistics with $f_{+}$. Surprisingly, this braiding statistics can also be (partially) detected in the physical Hilbert space, as we demonstrate below. First, note that $\rho_{f}$ preserves a weak $1$-form symmetry generated by $e$ anyons, $W^{e}_{\gamma}\rho_{f}W^{e}_{\gamma}=\rho$ for contractible closed $e$ string $W^{e}_{\gamma}$ de Groot et al. (2022). Then based on Definition 1, deconfined excitations can be created using an open $e$ string, $\rho_{f}\rightarrow\rho^{e}_{f}=W^{e}_{C}\rho_{f}W^{e}_{C}$. Secondly, introduce an ancilla qubit and prepare the initial state $|+\rangle\langle+|\otimes\rho_{f}^{e}$. Then use the ancilla qubit to control the braiding process: if the ancilla is in $|1\rangle$, we create two $f$ anyons and drag one of them along a loop $l$ enclosing the $e$ anyon, and finally annihilate with the other $f$ anyon; if the ancilla is in $|0\rangle$ we do nothing. This controlled process can be performed using the unitary gate $V=|0\rangle\langle 0|\otimes I+|1\rangle\langle 1|\otimes W^{f}_{\tilde{\gamma}}$, where $W^{f}_{\tilde{\gamma}}=\prod_{i\in\text{loop }\tilde{\gamma}}X_{i}Z_{i+\bm{\delta}}$. Then the braiding statistics can be detected by rotation of the ancilla qubit:
@@
One key distinction to the usual anyon braiding is that the $f-e$ braiding here is only one-way defined under the above protocol. Namely, the statistical phase can only be detected by moving $f$ around $e$ (with the movement of $f$ controlled by the ancilla), but not the other way around  [^2] Nevertheless, one can alternatively use the ancilla to control the creation of two $f$ anyons and extract the braiding statistics by moving $e$ around $f$.. This directly follows from the fact that while $f$ anyons generate a strong $1$-form symmetry, $e$ anyons only generate a weak $1$-form symmetry. Therefore, a finer characterization of deconfined excitations than Definition 1 is needed, and we call the former type as strongly deconfined and the latter type as weakly deconfined. As demonstrated above, strongly deconfined anyons are allowed to form coherent superposition states, and their statistics can be detected from the interference effects. Weakly deconfined anyons, on the other hand, always have trivial exchange and braiding statistics among themselves and can have one-way braiding statistics with strongly deconfined anyons. In the rest of the paper we mainly focus on strongly deconfined anyons.
@@
Using the perspective of anomalies, we can generalize our construction to obtain other intrinsic mixed-state TO Sohal and Prem (2024); Ellison and Cheng (2024). Given a TO with anyon content $\mathcal{A}$, one can incoherently proliferate a subset $\mathcal{B}\subset\mathcal{A}$ of anyons using a noisy channel, taking the Kraus operator to be the shortest string of anyons in $\mathcal{B}$. If there exists some Abelian anyon $a\in\mathcal{A}$ that have trivial braiding with all anyons in $\mathcal{B}$, then closed $a$ strings commute with the Kraus operators. Therefore, the $a$ anyons generate a strong $1$-form symmetry Li and Mong (2024). Based on the previous discussion, it means they remain deconfined excitations with well-defined statistics. When they have nontrivial statistics, $\theta(a)\neq 1$, then the 1-form symmetry is anomalous. For example, in the model discussed above, $\mathcal{A}=\{1,e,m,f\},\mathcal{B}=\{f\}$, and $a=f$ because fermions have trivial full braiding with themselves.
The existence of non-bosonic strongly deconfined excitation can be viewed as another diagnosis of nontrivial mixed-state TO, in complementary to TEN. Indeed, we show below that this property generically implies the mixed state to be LRE, without explicit calculation of negativity.
@@
    Typically, the 1-form symmetry includes generators supported on non-contractible loops. However, here we only require the symmetry condition for contractible ones, enabling a much wider range of applications. For example, for the construction of intrinsic mixed-state TO discussed above, the pre-decoherence state can be taken to be any state in the ground state subspace. The above theorem can also be applied to topologically trivial spatial manifolds like a $2$-sphere.
@@
    To prove the theorem, we make an additional assumption that $W^{a}_{\gamma}$ can be expressed using an FDLUC, which is usually the case for Abelian anyon $a$.
Proof. Suppose by contradiction that $\rho=\mathcal{N}_{\text{FDLC}}[\sum_{\lambda}p_{\lambda}\bigotimes_{\text{site}i}|\psi^{\lambda}_{i}\rangle\langle\psi^{\lambda}_{i}|]$. Recall that a generic FDLC $\mathcal{N}$ can be constructed in three steps: First, introduce auxiliary degrees of freedom on each site. Secondly, apply a finite-depth local unitary circuit (FDLUC) on the system and the auxiliary degrees of freedom. Finally, trace out the added degrees of freedom. It can be written as $\mathcal{N}_{\text{FDLC}}[\rho_{0}]=\text{tr}_{E}[U_{SE}\rho_{0}\bigotimes_{\text{site i}}|e_{i}\rangle\langle e_{i}|U^{\dagger}_{SE}]$, where $U_{SE}$ is a FDLUC. If $\rho_{0}$ is a fully separable state, then $\rho=\mathcal{N}_{\text{FDLC}}[\rho_{0}]$ is of the form $\rho=\sum_{\lambda}p_{\lambda}\text{tr}_{E}[U_{SE}\bigotimes_{i}|\phi_{i}^{\lambda}\rangle\langle\phi_{i}^{\lambda}|U^{\dagger}_{SE}]$, where $|\phi_{i}\rangle\equiv|\psi_{i}^{\lambda}\rangle\otimes|e_{i}^{\lambda}\rangle$ Then each pure state $U_{SE}\bigotimes_{i}|{\phi}_{i}\rangle$ must be symmetric: $W^{a}_{\gamma}U_{SE}\bigotimes_{i}|\phi_{i}\rangle=U_{SE}\bigotimes_{i}|\phi_{i}\rangle$. Equivalently, $\tilde{W}^{a}_{\gamma}=U^{\dagger}_{SE}{W}_{\gamma}^{a}U_{SE}$ is a symmetry of the product state $\bigotimes_{i}|\phi_{i}\rangle$. Crucially, the restriction of $\tilde{W}^{a}$ to open strings $\tilde{W}^{a}_{i_{1}i_{2}}\equiv U^{\dagger}_{SE}{W}_{i_{1}i_{2}}^{a}U_{SE}$ satisfies the same algebra (18) as $W^{a}$. As we show below, this will lead to contradiction.
@@
By applying both sides of the equation on $\bigotimes_{i}|\phi_{i}\rangle$, we get $A_{k}\otimes B_{j}\otimes A_{l}\otimes B_{m}\bigotimes_{i}|\phi_{i}\rangle=\theta(a)A_{k}\otimes B_{m}\otimes A_{l}\otimes B_{j}\bigotimes_{i}|\phi_{i}\rangle$, which leads to contradiction when $\theta(a)\neq 1$.
The above proof shows that nontrivial TO is guaranteed by the existence of non-bosonic deconfined excitation $a$, even for mixed states. Furthermore, when $a$ is neither bosonic or fermionic, $\theta(a)\neq\pm 1,$ or when $a$ has nontrivial mutual statistics with another strongly deconfined anyon $a^{\prime}$, then an even stronger conclusion can be proved.
Theorem 2: Consider a 2-dimensional state $\rho$ with an anomalous strong 1-form symmetry generated by Abelian anyons $a$, $a^{\prime}$ with nontrivial braiding statistics $B_{\theta}(a,a^{\prime})\neq 1$ (with the special case $a=a^{\prime}$ included). Then for any bipartition $A\bigcup B$ with linear size $L_{A},L_{B}\rightarrow\infty$ in the thermodynamic limit, $\rho$ cannot be prepared using a FDLC from any bipartite separable state:
@@
$\tilde{W}^{a^{\prime}}_{j^{\prime}_{A}i^{\prime}_{A}}\tilde{W}^{a}_{j_{A}i_{A}}=B_{\theta}(a,a^{\prime})\tilde{W}^{a}_{j_{A}i_{A}}\tilde{W}^{a^{\prime}}_{j^{\prime}_{A}i^{\prime}_{A}}$ (25)
$\Rightarrow$ $u^{\dagger}_{\sigma_{i^{\prime}j^{\prime}}}\tilde{W}^{a^{\prime}}_{j^{\prime}_{A}i^{\prime}_{A}}u^{\dagger}_{\sigma_{ij}}\tilde{W}^{a}_{j_{A}i_{A}}=B_{\theta}(a,a^{\prime})u^{\dagger}_{\sigma_{ij}}\tilde{W}^{a}_{j_{A}i_{A}}u^{\dagger}_{\sigma_{i^{\prime}j^{\prime}}}\tilde{W}^{a^{\prime}}_{j^{\prime}_{A}i^{\prime}_{A}},$
@@
IV Generalizations to other intrinsic mixed-state TO